OpenStudy (anonymous):

What is the value of the expression? 4!(7 – 4)! A. 144 B. 120,384 C. 5040 D. 479,001,600

OpenStudy (anonymous):

I don't understand the question. Can you re-write it please.

OpenStudy (anonymous):

Thats how it is on my practice test . I dont understand it either thats why Im here but I guess you can take out the ----> "!" .

OpenStudy (anonymous):

So just this 4(7 - 4). thats easy. :P

OpenStudy (anonymous):

I guesss :/

OpenStudy (anonymous):

Just use the distributive property.

jimthompson5910 (jim_thompson5910):

no, 4! = 4*3*2*1 = 24

jimthompson5910 (jim_thompson5910):

and (7-4)! = 3! = 3*2*1 = 6

jimthompson5910 (jim_thompson5910):

So 4!(7 – 4)! = 24*6 = 144

OpenStudy (anonymous):

@jim_thompson5910 so its 144 ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

@jim_thompson5910 , thanks :)

jimthompson5910 (jim_thompson5910):

yw